TYPESETTING CAMERA-READY MANUSCRIPTS FOR LAKE LOUSE WINTER INSTITUTE PROCEEDINGS

F.C. KHANNA

Centre for Subatomic Research, Department of Physics, University of Alberta, Edmonton, Alberta, Canada T6G 2N5

ABSTRACT

Use upper case letters for the title and your name. Use upper and lower case letters for your address. If there are multiple authors with one address they can be on one author line. If there are multiple authors with multiple addresses, the sequence \author{} – \address{} may be repeated any number of times. You may use \\ any place in the title, author, or address in order to force good line breaks.

1 Sectioning and Referencing

1.1 Sectioning

The \section{} and \subsection{} commands will provide the correct heading styles if you use upper and lower case letters for the headings. Use the form “Section 2.2” when referring to a section in the text. The command \label{key} immediately following the \(sub)section{} command will assign the number of the (sub)section to the label “key” and you can refer to it using \ref{key}.

1.2 Referencing

Citations are handled with the LaTEX[1] command \cite{key}. Please insure that the order of items in the “Reference” section at the end of the text is the same as the order that the items appear in the main text. References should provide initials and last name of the author(s), title of the publication (italics), volume (boldface), year of publication in the journal/book and the page number. Some macros to simplify journal citations are included near the beginning of this file. Here is an example[2,3] of their use.

2 Tables and Figures

2.1 Tables

You should use the \table{} environment for tables. This insures that the captions are correctly placed set. Table 1 is an example of a table. Notice that there is a single line bordering the table on all sides and a single line beneath the column headers. You can use vertical lines to separate columns, but some people think that tables look better without them.

Table 1: This is a simple table but it illustrates how to make a table and to typeset a caption for it, as well as generate a label for reference to the table.

Variable Value
a  0
b  1

You should locate a table as close to the relevant text as possible. The optional argument [htb] tells LaTEX to first try to place the table in the current location (h) then at the top of the page (t) and then the bottom of the page (b). You can choose a different order if you prefer. Make sure that the table fits within the boundaries of the text (8.5 × 6.0 in) and that the table is centered as illustrated.

2.2 Figures

You should use the \figure{} environment for figures. Figures are handled in the same way as tables, but the caption is below the figure instead of above it, as indicated in Figure 1. Make sure that the figure and its caption fit within the boundaries of the text and that the figure is centered.

Figure 1: A space of only 2 cm has been left for inserting a figure using real cut and paste. If you insert an EPS figure directly, eliminate the \vspace∗{2cm} commands and use the sequence, \begin{center} … \end{center} instead. The … represent your favorite macro for installing an EPS figure.

3 Equations

Displayed equations should be created with the LaTEX equation or eqnarray environments which will automatically give consecutive numbers to the equations. Again you can use the \label{key} command in these environments to define a key for this equation which you can reference elsewhere in the text using \ref{key}, e.g., Eq. (1).

B(a → b) ≡ \frac{Γ(a → b)}{Γ}  (1)

Acknowledgments

Acknowledgments should be placed at the end of the text before references. There should not be a section number, so you should use \section∗{Acknowledgements}.

Do the same with any appendices and give them section names “Appendix A”, “Appendix B”, etc., by hand.

References

1. L. Lamport, LaTEX: A Document Preparation System (Addison-Wesley, Reading, 1968)

2. A.B. Author et al., Phys. Rev. Lett. 1, 123 (1990)

3. B.C. Author, Phys. Lett. B2, 234 (1980)